Maria drove from Chicago, Illinois, to Milwaukee, Wisconsin, a distance of 90 miles, at a mean speed of 55 miles per hour. On her return trip, the traffic was much heavier, and her mean speed was 40 miles per hour. Find Maria's mean speed for the round trip.

Answer by mananth(16946)

total distance of the trip / total time up & down
..
time while going to milwaukee = 90/55 = 1.6 hours
time on return journey = 90/40= 2.25 hours
Total time = 2.25+1.6= 3.85 hours
...
round trip distance = 180 miles
Average speed = 180 / 3.85 = 46.75 mph
...
m.ananth@hotmail.ca